### v1

@@ Abstract
Priyanshu Kumar [^1] Equal contributors, correspondence at {priyansk, devanshj}@cs.cmu.edu Devansh Jain [^1] Equal contributors, correspondence at {priyansk, devanshj}@cs.cmu.edu Akhila Yerukola
@@ 1 Introduction
Overall, our datasets and models [^2] Model, code, and data will be made public upon acceptance under the ODC-BY license. serve as a starting point for building powerful and robust multilingual safety detectors and advance efforts towards multilingual safe AI systems.
@@ 2.4 Human Validation
To validate the translation quality and the generated safety labels, we conduct human validation across all 16 languages. Due to budget constraints, we randomly sample 50 data points per language, ensuring a balanced distribution across PGMix (train) and PGPrompts (test), harmful and harmless labels, as well as user prompts and model responses. We recruit workers from Prolific [^6] https://www.prolific.com, filtering them based on their proficiency in each language. Each data point is evaluated by three annotators.
@@
    Translation Quality: Using the Direct Assessment + Scalar Quality Metric (DA+SQM) framework Kocmi et al. (2022), we elicit a score between 0 and 100 on a continuous sliding scale with seven labeled tick marks.
@@
Figure 4: Safety category distributions for PGMix ITW samples. Takeaway: ITW and WildGuardMix have different safety category distributions.
@@ 3 PolyGuard: A 17-Language Safety Moderation Tool
To build PolyGuard, we fine-tune Qwen2.5-7B-Instruct (Yang et al., 2024a) and Ministral-8B-Instruct-2410, both of which have been shown to have state-of-the-art performance in multilingual knowledge and commonsense, code, and math settings Qwen ; Mistral . We refer to these models as PG Qwen2.5 and PG Ministral In addition, we also fine-tune Qwen2.5-0.5B-Instruct to build PG Smol.
@@ 4 Results & Research Questions
We compare PolyGuard with popular open-source safety detection models of similar size (Yang et al., 2024b), namely Llama-Guard-2 (Team, 2024), Llama-Guard-3-8B (Dubey et al., 2024), Aegis 1.0 Defensive (Ghosh et al., 2024), MD Judge (Li et al., 2024), and DuoGuard Deng et al. (2025). We also benchmark proprietary models, namely Perspective API [^7] https://perspectiveapi.com/, OpenAI Omni Moderation [^8] https://platform.openai.com/docs/models/omni-moderation-latest, and Google Moderation [^9] https://cloud.google.com/natural-language/docs/moderating-text.
@@
Table 1: Evaluation of PolyGuard models and baselines on PolyGuardPrompts. Takeaway: Both PG models perform equally well on in-distribution data and outperform baselines.
@@ 4.1 How do PG models perform on the in-distribution PGPrompts benchmark?
PG models based on Qwen2.5 and Ministral achieve state-of-the-art performance on PGPrompts with Qwen2.5 performing marginally better. PG Smol outperforms DuoGuard, its similar size counterpart (Table 1). Aegis Defensive supports only a single text as input and is hence evaluated for Harmful Request only. Since the remaining baselines do not explicitly support Harmful Response, we approximate the prediction by executing them on prompt + response. None of the baselines support the Response Refusal task. Out of all baselines, the safety category taxonomy is the same for Llama-Guard-3 and PG. We observe that Llama-Guard-3 achieves marginally better performance for Response Safety Violations task because it conservatively predicts only safety category for most of the samples in PGPrompts; PG, on the other hand, predicts multiple violations, thus leading to lower Exact Match and comparable Jaccard similarity scores.
@@ 4.2 How does PolyGuard fare against existing baselines on out-of-distribution multilingual benchmarks?
Table 2: F1 scores of safety detectors on Multilingual Guardrail Test Suite; metrics are in bold and underlined for the best second-best performing models respectively. Mod.=Moderation, XS=XSafety, MJ=MultiJail, En.=English, Mul.=Multilingual, LG=Llama Guard.
@@
Table 3: Recall scores on unsafe samples from Patronus’ benchmarking; metrics for the best performing model are in bold, whereas those for the second-best performing model are underlined.
@@ 4.4 How does performance vary on English vs Translated vs Code-Switched data?
We evaluate models on the Code-Switching Red-Teaming (CSRT) Yoo et al. (2024) dataset and the translated and code-switched version of Aegis 1.0 (Ghosh et al., 2024) as provided by Yang et al. (2024b). Since CSRT also evaluates LLMs’ tendency to generate unsafe content, we use the same automatic annotation pipeline as described in Appendix D.
@@ 5 Background & Related Work
AI Safety, the field of research focused on ensuring that AI systems are developed and deployed in a manner that is trustworthy, responsible, reliable, and beneficial to humans Chen et al. (2024), has become widely studied in recent years (Chua et al., 2024; Hendrycks, 2025; Bengio et al., 2025; Bullwinkel et al., 2025). This increasing interest has led to the procurement of datasets for training and evaluating safety guardrails for AI systems (Ghosh et al., 2024; ; Han et al., 2024; Lin et al., 2023; Ji et al., 2023; Li et al., 2024). Similarly, safety benchmarks have been curated to evaluate the safety risks exhibited by AI systems (Xie et al., 2024; Mazeika et al., 2024; Jain et al., 2024; Kumar et al., 2024; Yoo et al., 2024; Zeng et al., 2024b; Zhang et al., 2024a; b; Tan et al., 2024). However, almost all of the aforementioned datasets are limited to the English or Chinese language only or focus on specific subsets of AI safety Jain et al. (2024).
@@ Acknowledgments
This research was supported in part by Jigsaw, DSO National Laboratories and Microsoft’s Accelerating Foundation Models Research program.
@@ Appendix A Translation Pipeline and Quality Evaluation
Tables 6 and 7 show the verdicts of the GPT-4o judge for the human prompt and model response respectively. We observe that TowerInstruct generates higher-quality translations when compared to NLLB for the languages it supports. However, in the case of Hindi (which is not supported by Tower), the quality is poor.
@@ Appendix B Human Validation
We collect 2,400 annotations across 16 languages and 50 data points per language, with each data point annotated by 3 annotators, and each annotator annotating 10 data points. We recruited 191 unique annotators141414some participated in multiple languages, resulting in a lower unique count via Prolific, spanning across 24 countries. They self-identified as 110 male and 81 female. In terms of ethnicity, they described themselves as 84 White, 79 Black, 12 Mixed, 10 Asian, and 5 Other.
@@ Appendix C PolyGuard Training Details
We train our models using openrlhf [^15] https://github.com/OpenRLHF/OpenRLHF/tree/main on 8 NVIDIA A6000 GPUs. We set LoRA rank to 8 and alpha to 16. We train our models with a total batch size of 128, for a sequence length of 8192, for 1 epoch using a learning rate of $42e-$ 4. The system and user prompts (adapted from WildGuard and Llama Guard v3) used by PG are as follows:
@@ Appendix F Influence of low-quality translated data
Intuitively, the presence of poor-quality translated data should harm model performance. However, PG models show contrastive trends when low-quality samples are removed from the training data mix (Figure 9). The performance of Qwen2.5 degrades for most datasets, whereas the performance of Ministral improves. The performance degradation in the case of Qwen2.5 can be attributed to noisy samples in safety and toxicity evaluation datasets. Harmful text is considered to belong to low-quality data; web-crawls implement word blocklist filters to enhance data quality Dodge et al. (2021). Thus, we hypothesize that the noise induced by poor translations bridges the gap between training and evaluation data, thus leading to performance improvement.

### v2

@@ Abstract
Priyanshu Kumar [^1] Equal contributors, correspondence at msap2@cs.cmu.edu. Devansh Jain [^1] Equal contributors, correspondence at msap2@cs.cmu.edu. Akhila Yerukola
@@ 1 Introduction
Overall, our datasets and models [^2] Model, code, and data are available under the ODC-BY license. serve as a starting point for building powerful and robust multilingual safety detectors and advance efforts towards multilingual safe AI systems.
@@ 2.4 Human Validation
To validate the translation quality and the generated safety labels, we conduct human validation across all 16 languages. Due to budget constraints, we randomly sample 50 data points per language, ensuring a balanced distribution across PGMix (train) and PGPrompts (test), harmful and harmless labels, as well as user prompts and model responses. We recruit workers from Prolific, [^6] https://www.prolific.com filtering them based on their proficiency in each language. Each data point is evaluated by three annotators.
@@
    Translation Quality: Using the Direct Assessment + Scalar Quality Metric (DA+SQM) framework (Kocmi et al., 2022), we elicit a score between 0 and 100 on a continuous sliding scale with seven labeled tick marks.
@@
Figure 4: Safety category distributions for PGMix ITW samples.
@@ 3 PolyGuard: A 17-Language Safety Moderation Tool
To build PolyGuard, we fine-tune Qwen2.5-7B-Instruct (Yang et al., 2024a) and Ministral-8B-Instruct-2410, both of which have been shown to have state-of-the-art performance in multilingual knowledge and commonsense, code, and math settings (Qwen, ; Mistral, ). We refer to these models as PG Qwen2.5 and PG Ministral In addition, we also fine-tune Qwen2.5-0.5B-Instruct to build PG Smol.
@@ 4 Results & Research Questions
We compare PolyGuard with popular open-source safety detection models of similar size (Yang et al., 2024b), namely Llama-Guard-2 (Team, 2024), Llama-Guard-3-8B (Dubey et al., 2024), Aegis 1.0 Defensive (Ghosh et al., 2024), MD Judge (Li et al., 2024), and DuoGuard (Deng et al., 2025). We also benchmark proprietary models, namely Perspective API [^7] https://perspectiveapi.com/, OpenAI Omni Moderation [^8] https://platform.openai.com/docs/models/omni-moderation-latest, and Google Moderation [^9] https://cloud.google.com/natural-language/docs/moderating-text.
@@
Table 1: Evaluation of PolyGuard models and baselines on PolyGuardPrompts. Takeaway: PG models outperform baselines on in-distribution data.
@@ 4.1 How do PG models perform on the in-distribution PGPrompts benchmark?
PG models based on Qwen2.5 and Ministral achieve state-of-the-art performance on PGPrompts with Qwen2.5 performing marginally better. PG Smol outperforms DuoGuard, its similar size counterpart (Table 1). Aegis Defensive supports only a single text as input and is hence evaluated for Harmful Request only. Since the remaining baselines do not explicitly support Harmful Response, we approximate the prediction by executing them on prompt + response. None of the baselines support the Response Refusal task. Out of all baselines, the safety category taxonomy is the same for Llama-Guard-3 and PG. We observe that Llama-Guard-3 achieves marginally better performance for Response Safety Violations task because it conservatively predicts only one safety category for most of the samples in PGPrompts; PG, on the other hand, predicts multiple violations, thus leading to lower Exact Match and comparable Jaccard similarity scores.
@@ 4.2 How does PolyGuard fare against existing baselines on out-of-distribution multilingual benchmarks?
Table 2: F1 scores of safety detectors on Multilingual Guardrail Test Suite; metrics are in bold and underlined for the best second-best performing models respectively. Mod.=Moderation, XS=XSafety, MJ=MultiJail, En.=English, Mul.=Multilingual, LG=Llama Guard. Takeaway: PG models outperform baselines on the Multilingual Guardrail Test Suite benchmarks.
Multilingual Bench:
Patronus AI Bench:
@@
Table 3: Recall scores on unsafe samples from Patronus’ benchmarking; metrics for the best performing model are in bold, whereas those for the second-best performing model are underlined. Takeaway: PG models outperform baselines on Patronus AI’s benchmarks.
@@ 4.4 How does performance vary on English vs Translated vs Code-Switched data?
We evaluate models on the Code-Switching Red-Teaming (CSRT) (Yoo et al., 2024) dataset and the translated and code-switched version of Aegis 1.0 (Ghosh et al., 2024) as provided by Yang et al. (2024b). Since CSRT also evaluates LLMs’ tendency to generate unsafe content, we use the same automatic annotation pipeline as described in Appendix D.
@@ 4.5 How is performance affected by removing low-quality translated data?
4.6 Does PolyGuard superficially align with artifacts of machine-translated text only?
The use of machine-translated data for training PolyGuard models can lead to the hypothesis that models learn only to rely on machine-translation artifacts in the data to evaluate safety. To investigate if this behavior exists, we evaluate our models on the Aya Red-teaming dataset (Ahmadian et al., 2024), which consists of manually created 7,419 samples in 8 languages, thus lacking the noise patterns present in machine-translated texts. We do not observe empirical evidence supporting the hypothesis (Table 6). [^11] We also use the Aya Red-teaming dataset to assess the need for multilingual safety classifiers by translating it to English via TowerInstruct-7B-v0.2 and then evaluating an English-only classifier (Llama-Guard-3-8B). PG Qwen2.5 significantly outperforms this setup – achieving a higher recall in French (0.916 vs. 0.706), Russian (0.926 vs. 0.669) and Spanish (0.952 vs. 0.681) – highlighting the limitations of relying solely on translation for multilingual safety moderation.
Model Average Std Dev
PolyGuard Qwen2.5 87.01 8.27
PolyGuard Ministral 84.04 12.25
PolyGuard Smol 65.25 25.02
Table 6: Recall scores for PolyGuard models on human-written samples from the Aya RedTeam benchmark. Takeaway: PolyGuard models generalize on data from different distributions despite being trained only on machine-translated data.
5 PolyGuard Runtime Comparison
We have trained and open-sourced models of three sizes (0.5B, 7B, and 8B). While all three can run on consumer hardware, the 0.5B can benefit on-device or latency-critical applications. We also test the latency of our models on 7419 samples from the Aya RedTeaming dataset (Ahmadian et al., 2024) on an NVIDIA L40S GPU using VLLM (Table 7), and find that our 0.5B model has a high throughput. However, our 7B and 8B models run comparatively slower than their similarly sized Llama Guard counterparts. Compared to Llama Guard, PolyGuard models solve more tasks, and thus require longer prompts and generate more output tokens, which leads to increased runtime.
Model Size Input Tokens Output Tokens Time (m:ss)
Llama Guard 2 8B 1575800 27536 2:13
Llama Guard 3 8B 1657409 36364 2:14
PolyGuard Smol 0.5B 1870206 239337 0:31
PolyGuard Qwen2.5 7B 1870206 243043 3:27
PolyGuard Ministral 8B 1881052 242426 3:58
Table 7: Latency comparison of PolyGuard models on Aya RedTeaming Takeaway: Smol is highly efficient, whereas Qwen and Ministral are slower than LlamaGuards as PolyGuard models solve multiple tasks.
@@ 6 Background & Related Work
AI Safety, the field of research focused on ensuring that AI systems are developed and deployed in a manner that is trustworthy, responsible, reliable, and beneficial to humans (Chen et al., 2024), has become widely studied in recent years (Chua et al., 2024; Hendrycks, 2025; Bengio et al., 2025; Bullwinkel et al., 2025). This increasing interest has led to the procurement of datasets for training and evaluating safety guardrails for AI systems (Ghosh et al., 2024; ; Han et al., 2024; Lin et al., 2023; Ji et al., 2023; Li et al., 2024). Similarly, safety benchmarks have been curated to evaluate the safety risks exhibited by AI systems (Xie et al., 2024; Mazeika et al., 2024; Jain et al., 2024; Kumar et al., 2024; Yoo et al., 2024; Zeng et al., 2024b; Zhang et al., 2024a; b; Tan et al., 2024). However, almost all of the aforementioned datasets are limited to the English or Chinese language only or focus on specific subsets of AI safety Jain et al. (2024).
@@ Acknowledgments
This research was supported in part by Google Jigsaw, DSO National Laboratories and Microsoft’s Accelerating Foundation Models Research program.
@@ Appendix A Translation Pipeline and Quality Evaluation
Tables 8 and 9 show the verdicts of the GPT-4o judge for the human prompt and model response respectively. We observe that TowerInstruct generates higher-quality translations when compared to NLLB for the languages it supports. However, in the case of Hindi (which is not supported by Tower), the quality is poor.
@@ Appendix B Human Validation
We collect 2,400 annotations across 16 languages and 50 data points per language, with each data point annotated by 3 annotators, and each annotator annotating 10 data points. We recruited 191 unique annotators151515some participated in multiple languages, resulting in a lower unique count via Prolific, spanning across 24 countries. They self-identified as 110 male and 81 female. In terms of ethnicity, they described themselves as 84 White, 79 Black, 12 Mixed, 10 Asian, and 5 Other.
@@ Appendix C PolyGuard Training Details
We train our models using openrlhf [^16] https://github.com/OpenRLHF/OpenRLHF/tree/main on 8 NVIDIA A6000 GPUs. We set LoRA rank to 8 and alpha to 16. We train our models with a total batch size of 128, for a sequence length of 8192, for 1 epoch using a learning rate of 2e42e-4. The system and user prompts (adapted from WildGuard and Llama Guard v3) used by PG are as follows:
@@ Appendix F Influence of low-quality translated data
Intuitively, the presence of poor-quality translated data should harm model performance. However, PG models show contrastive trends when low-quality samples are removed from the training data mix (Figure 9). The performance of Qwen2.5 degrades for most datasets, whereas the performance of Ministral improves. The performance degradation in the case of Qwen2.5 can be attributed to noisy samples in safety and toxicity evaluation datasets. Harmful text is considered to belong to low-quality data; web-crawls implement word blocklist filters to enhance data quality (Dodge et al., 2021). Thus, we hypothesize that the noise induced by poor translations bridges the gap between training and evaluation data, thus leading to performance improvement.